OpenStudy (anonymous):

Find R:

OpenStudy (anonymous):

\[\sum_{k=1}^{\infty} 3r ^{k-1} =5\]

OpenStudy (anonymous):

@hartnn Could u possibly help with this?

hartnn (hartnn):

yes

OpenStudy (anonymous):

Ok great. How should I approach the problem?

hartnn (hartnn):

do you know the formula ? ..... wait , what is the topic ? is this related to geometric sequences ?

OpenStudy (anonymous):

Yes it is

hartnn (hartnn):

cool, so can you find the terms ? 1st term, 2nd terms, .... ?

OpenStudy (anonymous):

Yes, the 1st term is 1, the 2nd term is 3, the 3rd term is 9

hartnn (hartnn):

where did the 'r' go ? :O

OpenStudy (anonymous):

See I dont think i'm doing it correctly :O

OpenStudy (anonymous):

3r^0, 3r^1, 3r^2, 3r^3

OpenStudy (anonymous):

I dont understand what to do with the r variable.

hartnn (hartnn):

ok, so find the common ratio now, know what that is ?

OpenStudy (anonymous):

would it be 3?

hartnn (hartnn):

no, see common ratio = 2nd term/1st term = 3rd term/2nd term =.... so on so, whats 3r/3 = ... ? 3r^2/3r = ... ?

OpenStudy (anonymous):

So just r?

hartnn (hartnn):

yes! thats your common ratio, now you are almost done, 3,3r,3r^2 .... is an infinite geometric series. and the sum is given by \(\Huge S_\infty=\dfrac{a}{1-r} \) where a is 1st term and r is common ratio so your sum will be ?

OpenStudy (anonymous):

Ok, im familiar with the sum formula. It would be.. no such value?

OpenStudy (anonymous):

1/1-1 = 1/0

OpenStudy (anonymous):

ohhhhh wait.

hartnn (hartnn):

what is 'no such value?' see, a = 3 right ? , r= r so, 3/(1-r) = 5 find r from here

OpenStudy (anonymous):

yeah. i was thinking r = 1. lol

OpenStudy (anonymous):

r = 2/5

hartnn (hartnn):

correct! :)

OpenStudy (anonymous):

Ok I see what you did there. Thanks so much. I now understand it. I was making it more complicated than what it actually was.

hartnn (hartnn):

ok, you're welcome ^_^